3) Solve the following initial value problem. ( 1; 0 <t y + y = f(t), y(0) = 2, y(0) = -1, where f(t) = } nere -1; En VI t
Answer #1

y + y = f(t), 7(0) = 2, 910=-1 where f(L) = { 1; OSE LA Here the auxilary equation is D²+1=0 Defi :. GF = A cosnt B sinn Dzi